# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

### v4

@@ 2.3 Pricing on Non-Convex Markets
Once an allocation $(x,y)$ has been obtained, the market operator must provide electricity prices $p\in\mathbb{R}^{N\times T}$ for each node and time period. Due to the non-convexities implied by sellers’ cost functions (i.e., the binary commitment variables $u_{st}$), this is not a trivial task. In microeconomics, the welfare theorems provide a foundation for pricing in markets. In their seminal paper, Arrow & Debreu, (1954) demonstrate that under convex preferences, demand independence, and perfect competition with divisible items, a market operator can always achieve a set of Walrasian equilibrium prices that support the welfare-maximizing allocation. We assume quasilinear utilities, i.e., the utility of each market participant is defined as the difference between valuation or costs and price.
@@
Unfortunately, real-world electricity markets do not satisfy the assumptions underlying the welfare theorems. As in many other markets, items are not perfectly divisible and market participants exhibit non-convex preferences. A large stream of literature has investigated existence conditions for Walrasian equilibria in non-convex markets (Kelso & Crawford,, 1982; Bikhchandani & Ostroy,, 2002; Baldwin & Klemperer,, 2019), but none of these conditions are satisfied in electricity markets. Bikhchandani & Mamer, (1997) demonstrate that Walrasian equilibria exist if and only if the linear relaxation of the allocation problem has an integer solution. When this is not the case, market participants bear (global) lost opportunity costs at any set of prices and do not maximize their individual profits, violating envy-freeness.
Consequently, market operators have to resort to heuristic pricing rules that aim at approximating Walrasian equilibrium prices. Many such pricing rules have been proposed (Liberopoulos & Andrianesis,, 2016), each compromising on the properties of Walrasian equilibria differently. In this work, we focus on Convex Hull (CH) pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) and Integer Programming (IP) pricing (O’Neill et al.,, 2005) as two established pricing rules in practice, as well as on the Join pricing rule (Ahunbay et al.,, 2024) as a novel multi-objective approach to pricing. We also consider the Euphemia pricing rule as it is applied in European electricity markets.
2.3.1 Convex Hull Pricing
In the absence of Walrasian equilibrium prices, there is no set of prices $p$ that maximizes the profit of every market participant. The concept of global lost opportunity costs describes this forgone payoff.
@@ 2.3.1 Convex Hull Pricing
The key idea of CH pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) is to minimize GLOCs over all market participants. Formally, CH pricing replaces each non-convex cost function $c_{s}(y_{s})$ with its convex envelope in (1), and derives prices from the dual of the resulting convex problem.
Computing convex envelopes and determining CH prices is generally computationally challenging (Schiro et al.,, 2016). However, it has been shown by Hua & Baldick, (2017) that CH pricing becomes tractable when the cost function is described as in (3). Specifically, by relaxing the binary constraints $u_{st}\in\{0,1\}$ to $u_{st}\in[0,1]$ and solving the dual of (1), CH prices can be obtained from a single linear program. This approach, also referred to as Extended Locational Marginal Pricing (ELMP), was also used by the recent ENTSO-E study (ENTSO-E,, 2022).
2.3.2 IP Pricing
In a convex market, Walrasian equilibrium prices are equivalent to marginal prices, i.e., the costs of the last accepted bid set the prices. Following this notion, O’Neill et al., (2005) try to generalize marginal pricing for non-convex markets. Their IP pricing rule assumes that generators cannot easily deviate from their commitment status (as encoded by the binary variable $u_{st}$) and sets prices equal to the variable costs of the marginal committed unit (i.e., the marginal unit among all units with $u_{st}=1$).
IP pricing has gained popularity in many practical markets. It involves three steps: (i) obtaining the optimal commitment variables $u_{st}^{*}$ from the allocation problem in (1), (ii) fixing the commitment variables of each generator to these optimal values, i.e., setting $u_{st}\in\{0,1\}$ to $u_{st}\in[0,1]$ with $u_{st}=u_{st}^{*}$ in (3), and (iii) solving (1) with these linearized cost functions and extracting prices from its dual.
@@ 2.3.2 IP Pricing
By convexity of $v_{b}(x_{b})$, a buyer’s LLOCs exactly equal their GLOCs. Besides computational tractability, IP prices accurately reflect the marginal value of transmission capacity (Yang et al.,, 2019). Specifically, price differences among nodes arise solely when the network experiences congestion. It has been shown that this property immediately corresponds minimizing LLOCs (Ahunbay et al.,, 2024), making them a significant indicator of good congestion signals.
2.3.3 Join Pricing
@@
Bids that require MWPs are also known as paradoxically accepted bids. MWPs can be regarded as another subset of GLOCs, referring only to participants’ incentives to not participate in the market at all. However, both CH and IP pricing can experience high levels of MWPs. Recent practical concerns have thus motivated the development of pricing rules that ensure lower levels of MWPs. Among them is the Join pricing rule (Ahunbay et al.,, 2024), which proposes a dual pricing problem that combines the objective of IP pricing (minimizing LLOCs) with minimizing MWPs. Specifically, for each participant it considers the maximum of LLOCs and MWPs in the objective, leading to a robust and computationally tractable pricing rule with adequate congestion signals and low MWPs simultaneously.
2.4 Euphemia
@@
The Euphemia algorithm combines market clearing and pricing (unlike the pricing rules above which can be applied to any pre-computed allocation). In European markets, it computes an allocation and zonal prices, but theoretically, it can also be applied in nodal markets. Algorithmically, it starts by computing the welfare-maximizing allocation, solving (1). It then advances to examining whether prices can be established for the allocation, i.e., they must prevent any paradoxically accepted bids, ensure the acceptance of all in-the-money hourly bids and, optionally, certify that cross-zonal flows occur from a lower-price to a higher-price zone. If paradoxically accepted bids exist, Euphemia adds cuts to (1) that eliminate the current candidate solution, constraining the allocation problem and leading to welfare losses. These iterations between allocation and pricing continue until a solution without paradoxically accepted bids is reached.
Subsequently, Euphemia continues by addressing some of the idiosyncrasies of the European market (PUN search, re-insertion of certain paradoxically rejected bids, resolving indeterminacy). If, at any of these steps, violations are detected, the iterative procedure continues. Eventually, Euphemia outputs allocation, zonal prices, and cross-zonal flows.
@@
A practical problem with Euphemia is its computational tractability. Due to its iterative nature, multiple mixed-integer problems must be solved until allocation and prices are obtained. With the planned 15-min market time unit introduction in Europe, the scalability of Euphemia poses a significant concern, and policymakers consider non-uniform pricing rules (such as the ones introduced above) to obtain solutions faster (All NEMO Committee,, 2022).
3 Data and Processing
This study examines welfare, prices, and MWPs for multiple zonal/nodal configurations, allocation rules, and pricing mechanisms. As discussed, we leverage the data released in the context of the European bidding zone review (ENTSO-E,, 2022; ACER, 2022b, ). We focus on Germany as arguably one of the most important bidding zones regarding clearing volume and its central position in the European grid. In this section, we provide an overview of the data set and our processing of it.
3.1 Bidding Zone Review Data
Let us discuss the data in more depth. The topology of the grid, defined by its nodes and connections, dictates the path electricity takes from generation to consumption, with substations facilitating the necessary voltage transformations along the way. Substations act as critical nodes in the network that manage voltage levels for efficient power flow. A substation might handle multiple voltage levels. For example, a transmission substation might receive power at a very high voltage (e.g., 220kV or higher) and step it down to a lower high voltage (e.g., 110kV) for further transmission or to a medium voltage (e.g., 10-35kV) for distribution to urban or industrial areas.
The ENTSO-E dataset provides a model of Germany’s power grid, detailing 834 substations. Each of these substations encompasses one or more voltage levels, summing up to 1697 voltage levels at substations in the dataset. These voltage levels cover the spectrum from transmission ($\geq$ 220kV) to distribution ($\leq$ 110kV) and are composed of several topological nodes. Altogether, there are 2898 topological nodes listed in the data sets. Additionally, the dataset includes details for 100 substations located in the neighboring countries, which have been integrated into our analysis. For our study, we manually determined the geographical locations of all substations and computed prices for all voltage levels on these substations. After some pre-processing, which involved removing disconnected nodes, breakers, and disconnectors, our German grid model consists of 1670 nodes and 2470 lines (Schmitt,, 2023). Technically, supply and demand on different voltage levels of substations do not necessarily need to be the same. However, it turned out that prices computed for voltage levels on a substation were identical or very close in the ENTSO-E dataset, which is also the case for the nodal prices provided by ENTSO-E. Overall, the average prices that we computed were not significantly different from those in the BZR dataset for substations. There were differences on individual nodes, which is probably due to the fact that the prices computed by ENTSO-E are based on the duals of a linear programming relaxation, while we use the mixed integer linear program and pricing rules as they are defined above.
As the maximum power of each seller is given as a fixed quantity (denoting its nominal capacity), we need to manually incorporate the variability of renewable energy sources. We thus distribute the aggregate dispatched energy in each hour equally to all plants of the same type, proportional to their nominal capacities, and set the maximum power to this value. The underlying assumptions are that renewables are always dispatched at maximum power and that meteorological conditions are equal across Germany. Because hydro power plants are dispatchable, they are excluded from this logic. The data also do not include information on variable or fixed costs for renewables, which we derived from literature (Kost et al.,, 2021; IRENA,, 2022; ENTSO-E,, 2022). For thermal units, variable cost information are provided, while fixed cost information were again obtained from literature (Kost et al.,, 2021; Schröder et al.,, 2013; Bundesnetzagentur,, 2022).
An additional challenge is the mapping of generators to corresponding network nodes. The grid data specify the broad category of generating units at each node (e.g., thermal, hydro, external injections) and their nominal capacities. While renewable resources could be mapped to the seller data by matching nominal capacities, the grid data lacks additional information regarding thermal plant types. We assigned thermal generators in the grid to their operational characteristics (such as minimum uptime constraints) (Schmitt,, 2023). Firstly, we match identifiers with the Kraftwerksliste (power plant list) provided by the Bundesnetzagentur to obtain each generator’s broad thermal generation type (e.g., lignite, gas, etc.). As nominal capacities do not match between grid and seller data, and seller data have a higher granularity (e.g., Gas CCGT, Gas OCGT, etc.), we solve an integer program to assign each entity in the grid model to exactly one corresponding seller of the same plant type, with the objective to minimize the aggregated differences in capacities and number of units per plant type. We provide this integer program and the outcome of the generation matching in A. While this assignment might not be exact, we argue that smaller mistakes in the assignment of generators to nodes have little impact on the more general questions raised in our paper that address the benefits of different zonal and nodal configurations. The data are accessible through the ENTSO-E website.
@@ 3.2 Experimental Design
We implemented the models in Python 3.8 using Gurobi to solve optimization problems. We conducted experiments for eight weeks in 2009, as suggested by the ENTSO-E report (ENTSO-E,, 2022), for the German bidding zone. To maintain computational feasibility of calculating hourly allocations and prices, we permitted a MIP gap of 5% for mixed-integer programs. Table 1 summarizes our experimental design and the scope of our analysis.
We first employ different allocation rules in accordance with the ENTSO-E report. These configurations range from a national single-zone model to a fully nodal model, as well as the ACER proposals for two to four bidding zones ACER, 2022b . Each allocation is associated with generation costs as well as redispatch costs to obtain a physically feasible outcome. After computing an allocation, different pricing rules, such as IP, CH, or Join pricing, can be applied. Each pricing rule implies different price levels as well as MWPs for market participants. Prices were capped at 100 EUR/MWh to limit the impact of outliers. Note that prices higher than 100 EUR/MWh are regarded as outliers also in the BZR Report (ENTSO-E,, 2022). We also consider the Euphemia allocation and prices as the current implementation in European markets.
@@ 4 Results
In this section, we summarize the main results of our analysis. We start with comparing average generation and total costs, before discussing price levels and MWPs. We report aggregate statistics for all weeks under consideration and further pick the days of February 18, March 10, and November 23 for an in-depth analysis of representative days with low, medium, and high prices, respectively.
@@ 4.1 Generation and Redispatch Costs
Table 2 includes an overview of the average daily generation and redispatch costs (both for cost-minimizing and volume-minimizing redispatch) for the allocation rules under consideration.
@@
Generation 33,507.00 33,702.87 33,697.10 33,578.77 33,717.41 36,150.7
Min-Cost Redispatch 3,567.84 3,421.81 3,457.25 3,546.4 3,449.39 0
Minimum Total Costs A 37,074.84 37,124.68 37,154.35 37,125.17 37,166.8 36,150.7
Min-Volume Redispatch 5,736.95 5,506.98 5,598.22 5,694.76 5,560.43 0
Minimum Total Costs B 39,243.95 39,209.85 39,295.32 39,273.53 39,277.84 36,150.7
@@
The numbers reported in Table 2 have realistic orders of magnitude as compared to those provided by the German Bundesnetzagentur (Bundesnetzagentur,, 2024). However, our goal is not to come up with a precise prediction of costs for 2025, but to obtain a relative comparison between different zonal configurations and nodal pricing. Adding network constraints cannot lead to lower objective function values in the cost minimization. As the national allocation rule neglects all transmission constraints, their generation costs are a lower bound for all other allocation rules. On average, the suggested zonal configurations require only slightly more generation costs, suggesting that adding few transmission constraints to the computation has little impact on the costs. In contrast, the nodal allocation rule includes the entire transmission network. Obviously, higher generation costs are required to reach an allocation with feasible power flows. On average, generation costs increase by 7.46% after including all network constraints for the nodal model compared to the national configuration before redispatch.
After redispatch, the national configuration with minimum cost redispatch has 2.5% higher costs than the nodal model, and 13.65% higher costs with minimum volume redispatch. Note that we assume that all generators can be redispatched in our model, such that redispatch costs in reality might be higher. In comparison, the average daily redispatch costs in Germany between 2020 and 2022 amount to roughly kEUR 7,200 (Bundesnetzagentur,, 2022). The average daily redispatch costs in Germany in 2022 were even at around kEUR 11,600, having increased by almost 100% compared to 2021. We assume zero redispatch for a nodal system that satisfies $\Psi^{DC}$, although in practice there will always be some minimal level of redispatch due to changing weather conditions and the simplified representation of network flows in the DCOPF. At first sight, it might be curious that redispatch costs can even increase with additional zones. B provides a simplified example to explain how this can happen.

### v6

@@ 2.1.3 Pricing on Non-Convex Markets
Once an allocation has been obtained, the market operator must provide electricity prices $p\in\mathbb{R}^{N\times T}$ for each node and time period. Due to the non-convexities implied by sellers’ cost functions (i.e., the binary commitment variables $u_{st}$), this is not a trivial task. In microeconomics, the welfare theorems provide a foundation for pricing in markets. In their seminal paper, Arrow & Debreu, (1954) demonstrate that under convex preferences, demand independence, and perfect competition with divisible items, a market operator can always achieve a set of Walrasian equilibrium prices that support the welfare-maximizing allocation. We assume quasilinear utilities, i.e., the utility of each market participant is defined as the difference between valuation or costs and price.
@@
Unfortunately, electricity markets do not satisfy the assumptions underlying the welfare theorems. As in many other markets, items are not perfectly divisible and market participants exhibit non-convex preferences. A large stream of literature has investigated existence conditions for Walrasian equilibria in non-convex markets (Kelso & Crawford, , 1982; Bikhchandani & Ostroy, , 2002; Baldwin & Klemperer, , 2019), but none of these conditions are satisfied in electricity markets. Bikhchandani & Mamer, (1997) demonstrate that Walrasian equilibria exist if and only if the linear relaxation of the allocation problem has an integer solution. When this is not the case, market participants bear lost opportunity costs at any set of prices, which violates envy-freeness.
Consequently, market operators have to resort to heuristic pricing rules that aim at approximating Walrasian equilibrium prices. Many such pricing rules have been proposed (Liberopoulos & Andrianesis, , 2016), each compromising on the properties of Walrasian equilibria differently. In this work, we focus on Convex Hull and Integer Programming pricing as two established pricing rules in practice, as well as on the Join pricing rule as a novel multi-objective approach. We also consider a simplified version of the Euphemia pricing rule as it is applied in Europe.
Convex Hull Pricing
In the absence of Walrasian equilibrium prices, there is no set of prices $p$ maximizing the profit of every market participant. Global lost opportunity costs describe this forgone payoff.
@@
The key idea of CH (Convex Hull) pricing (Hogan & Ring, , 2003; Gribik et al., , 2007) is to minimize total GLOCs. Formally, CH pricing replaces each non-convex cost function $c_{s}(y_{s})$ with its convex envelope in (1), and derives prices from the dual of the resulting convex problem.
Computing convex envelopes and determining CH prices is generally computationally challenging (Schiro et al., , 2016). However, it has been shown by Hua & Baldick, (2017) that CH pricing becomes tractable for the cost function we consider. Specifically, by relaxing the binary constraints $u_{st}\in\{0,1\}$ to $u_{st}\in[0,1]$ and solving the dual of (1), CH prices can be obtained from a single linear program. This approach, also referred to as Extended Locational Marginal Pricing (ELMP), was also used by the recent ENTSO-E study (ENTSO-E, , 2022).
IP Pricing
In a convex market, Walrasian equilibrium prices are equivalent to marginal prices, i.e., the costs of the last accepted bid set the prices. Following this notion, O’Neill et al., (2005) generalized marginal pricing for non-convex markets. Their IP (Integer Programming) pricing rule assumes that generators cannot easily deviate from their commitment status and sets prices equal to the variable costs of the marginal committed unit (i.e., the marginal unit among all units with $u_{st}=1$).
IP pricing involves three steps: (i) obtaining the optimal commitment variables $u_{st}^{*}$ from the allocation problem in (1), (ii) fixing the commitment variables of each generator to these optimal values, i.e., setting $u_{st}\in\{0,1\}$ to $u_{st}\in[0,1]$ with $u_{st}=u_{st}^{*}$, and (iii) solving (1) with these linearized cost functions and extracting prices from its dual.
@@
Besides computational tractability, IP prices accurately reflect the marginal value of transmission capacity (Yang et al., , 2019), i.e., price differences among nodes arise solely when the network experiences congestion. It has been shown that this property immediately corresponds minimizing LLOCs (Ahunbay et al., , 2024), making them a significant indicator of good congestion signals.
Join Pricing
@@
Accepted bids that require MWPs are known as paradoxically accepted bids. MWPs can be regarded as another subset of GLOCs referring to participants’ incentives to not participate in the market. Both CH and IP pricing can experience high levels of MWPs. Recent practical concerns have thus motivated the development of pricing rules that ensure lower MWPs. Among them is the Join pricing rule (Ahunbay et al., , 2024), which proposes a dual pricing problem that combines the objective of IP pricing (minimizing LLOCs) with minimizing MWPs. Specifically, for each participant it considers the maximum of LLOCs and MWPs in the objective, leading to a computationally tractable pricing rule with adequate congestion signals and low MWPs simultaneously.
In the U.S., MWPs are generally not covered directly by revenues obtained from the buyers through energy purchases in the day-ahead market but require additional funding mechanisms which depend on the ISO (Townsend et al., , 2014). [^14] MWPs are also referred to as bid productions cost guarantees (NYISO), revenue sufficiency guarantees (MISO) and operating reserve credits (PJM). ISOs often impose uplift charges or ancillary service fees on market participants to fund MWPs. These charges are ultimately passed on to electricity consumers, either directly through consumption fees or indirectly through higher electricity bills.
2.1.4 Euphemia
@@
The Euphemia algorithm combines market clearing and pricing (unlike the pricing rules above which can be applied to any pre-computed allocation). In European markets, it computes an allocation and zonal prices, but theoretically, it can also be applied in nodal markets. Algorithmically, it starts by computing the welfare-maximizing allocation, solving (1). It then advances to examining whether prices can be established for the allocation, i.e., they must prevent any paradoxically accepted bids, ensure the acceptance of all in-the-money hourly bids and, optionally, certify that cross-zonal flows occur from a lower-price to a higher-price zone. If paradoxically accepted bids exist, Euphemia adds cuts to (1) that eliminate the current candidate solution, constraining the allocation problem and leading to welfare losses. These iterations between allocation and pricing continue until a solution without paradoxically accepted bids is reached. Subsequently, Euphemia continues by addressing some of the idiosyncrasies of the European market (PUN search, re-insertion of certain paradoxically rejected bids, resolving indeterminacy). If, at any of these steps, violations are detected, the iterative procedure continues. Eventually, Euphemia outputs allocation, zonal prices, and cross-zonal flows.
@@
A practical problem with Euphemia is computational tractability. Due to its iterative nature, multiple MILPs must be solved until allocation and prices are obtained. With the planned 15-min market time unit introduction in Europe, the scalability of Euphemia poses a significant concern, and policymakers consider non-uniform pricing rules (such as the ones introduced above) to obtain solutions faster (All NEMO Committee, , 2022).
@@ 2.2 Experimental Design
We leverage the data released in the context of the European bidding zone review (ENTSO-E, , 2022; ACER, 2022c, ). We focus on Germany as arguably one of the most important bidding zones regarding clearing volume and its central position in the European grid. The data we use in our study and the processing steps are described in B.
We implemented the models in Python 3.11 using Gurobi to solve optimization problems. We conducted experiments for eight weeks in 2009 for the German BZ. To maintain computational feasibility of calculating hourly allocations and prices, we permitted a MIP gap of 5% for mixed-integer programs. Table 1 summarizes our experimental design and the scope of our analysis.
We first employ different allocation rules in accordance with the ENTSO-E report. These configurations range from a national single-zone model to a fully nodal model, as well as the ACER proposals for two to four bidding zones ACER, 2022c . Each allocation is associated with generation costs as well as redispatch costs to obtain a physically feasible outcome. After computing an allocation, different pricing rules, such as IP, CH, or Join pricing, can be applied. Each pricing rule implies different price levels as well as MWPs for market participants. Prices were capped at 100 EUR/MWh to limit the impact of outliers. Note that prices higher than 100 EUR/MWh are regarded as outliers also in the BZR Report (Section 2.4) (ENTSO-E, , 2022). We also consider the Euphemia allocation and prices as the current implementation in European markets.
@@ 3 Results
In this section, we summarize the main results of our analysis. We start with comparing average generation and total costs, before discussing price levels and MWPs. We report aggregate statistics for all weeks under consideration. C includes an analysis of representative days with low, medium, and high prices, respectively.
@@ 3.1 Generation and Redispatch Costs
Table 2 presents an overview of the average daily generation and redispatch costs for the allocation rules under consideration. Table 3 shows the percentage decrease of the nodal costs relative to the total costs associated with different zonal configurations. [^15] The values are obtained by computing $\frac{x\cdot 100}{y}-100$, where $x$ is the zonal cost for a specific configuration and $y$ is the nodal cost.
@@
Generation 32,145.08 32,428.55 32,219.17 32,198.95 32,382.74 36,150.7
RD-Min-Cost 6,140.06 5,687.83 5,863.74 5,827.86 5,726.9 0
Total RD-Min-Cost 38,285.14 38,116.38 38,082.91 38,026.81 38,109.64 36,150.7
@@
% National 2 Zones (k) 2 Zones (s) 3 Zones 4 Zones
Total RD-Min-Cost 5.90 5.44 5.34 5.19 5.42
Table 3: Percentage decrease of the nodal costs relative to the total costs associated with different zonal configurations
The numbers reported in Table 2 have realistic orders of magnitude as compared to those provided by the German Bundesnetzagentur (Bundesnetzagentur, , 2024). However, our goal is not to come up with a precise prediction of costs for 2025, but to obtain a relative comparison between different zonal configurations and nodal pricing. Adding network constraints cannot lead to lower objective function values in the cost minimization. As the national allocation rule neglects all transmission constraints, its generation costs are a lower bound for all other allocation rules. On average, the suggested zonal configurations require only slightly more generation costs, suggesting that adding few transmission constraints to the computation has little impact on the costs. In contrast, the nodal allocation rule includes the entire transmission network. Obviously, higher generation costs are required to reach an allocation with feasible power flows. On average, generation costs increase by 5.90% after including all network constraints for the nodal model compared to the national configuration before redispatch.
After redispatch, the national configuration has 5.90% higher costs than the nodal model. Note that we assume that all generators can be redispatched in our model, such that redispatch costs in reality might be higher. Also, in our zonal market clearing problem, we do not consider critical elements within zones and assume that the demand is inelastic and must be fully satisfied in any feasible allocation. In comparison, the average daily redispatch costs in Germany between 2020 and 2022 amount to roughly kEUR 7,200 (Bundesnetzagentur, , 2022). [^16] These costs include power- and voltage-related measures, countertrading and redispatch test procedures. The average daily redispatch costs in Germany in 2022 were even at around kEUR 11,600, having increased by almost 100% compared to 2021. We assume zero redispatch for a nodal system that satisfies $\Psi^{DC}$, although in practice a post-processing step is required to ensure that the DCOPF solution is ACOPF-feasible (Taheri & Molzahn, , 2024). In the US, this post-processing depends on the ISO and involves using an ACOPF model to adjust the DCOPF solution, accounting for real power losses, voltage magnitudes, and reactive power needs (California ISO, , 2024).